1. General status updates and schedule review
Glenn -- thinking about rack monitor box address assignment, and how we can use SD card. Going to reach out to Jason for testing at the test stand.

Tomo -- working on CRV data format running with package compatibility -- targeting collaboration meeting

Dariush and Stefano -- want to flesh out Calorimeter ROC details and slow controls PVs. Franco is coming soon. Dariush here for two weeks.

Greg -- integration team is happy with sync approach. Thinking about verifying chain CFO-DTCS[0:9]-ROC.. We want timestamping kits in DAQ room. And timestamping in DTC. Could take emulator ROC data in run data at something like 1 Hz. Repositories created for all detectors.

Jose -- place and routing now. Would like simulation help on what the DTC will send with new DCS format (single and double read, save block read for later).

Rick -- STM firmware ready. DCS changes introduced a bug into DTC - looking at.. John is getting a lot of quotes for fiber plant. Draka fiber order is on hold.. Gary is working on green light questions. STM limiting factor is we dont have timing cards.

Eric -- software updates for base packages.. artdaq and ots new releases out. Bug in artdaq release (that does not effect Mu2e).. Should be a new release next week or so. Req for 5 servers seems to be stalled.

2. Timing distribution demonstration updates
3. online Trigger Processing (including through DTC) progress/plans
4. Status of tests with detector subsystems
5. Rack protection & slow controls progress/plans (dedicated DCS node at FCC)